negative charge of -5.0x10^-4 C exerts an attractive force of 12 N on a second charge that is 0.040 m away. what is the magnitude of the second charge?


1
Expert's answer
2021-05-27T17:48:17-0400
"F=k\frac{q_1q_2}{r^2}\\12=9\cdot10^9\frac{5\cdot10^{-4}q_2}{0.04^2}\\q_2=4.3\cdot10^{-9}\ C"